How many grams of water are produced by burning 96 grams of H2 and the same amount of O2.


1
Expert's answer
2020-12-01T09:30:58-0500

2H2 + O2 = 2H2O

O2 is the limiting reagent

O2 = 2H2O

16×2 g O2 = 2 × 18 g H2O

1g O2 = 2×18 / 32 gH2O

96g O2 = 2×18 ×96 / 32 g H2O

= 108g H2O